# Finding product sets in some classes of amenable groups

### v1

@@ 2. Preliminaries
[Mac64, Theorem 1] Let $(X,\mu,T)$ be an ergodic $G$-system. Then there exist a compact group $K$, a closed subgroup $H$ of $K$ and a continuous group homomorphism $\alpha:G\to K$ with dense image such that the Kronecker factor of $(X,\mu,T)$ is measurably isomorphic to the $G$-system $(Z,m,R)$ given as follows: $Z=K/H$ is a homogeneous space, $m$ is the normalized Haar measure on $Z$ and $R=(R_{g})_{g\in G}$, where for each $g\in G$, $R_{g}:Z\to Z$ is given by $R_{g}(z)=\alpha(g)z$. The system $(Z,m,R)$ is called a rotation on the homogeneous space $Z$.
This proposition allows us to identify the Kronecker factor with a rotation on a homogeneous space, whenever $(X,\mu,T)$ is ergodic.
Characteristic factors for $G$-systems: The notion of characteristic factors will play a fundamental role later in one of our proofs. Characteristic factors are relevant to the ergodic averages in question, in each different problem. Here we have the following theorem for the characteristic factors with respect to some double averages that will concern us.
@@
$\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}f_{1}\otimes T_{g}f_{2}=\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}\>|\>Z)\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z)$
in $L^{2}(X\times X,\mu\times\mu)$. We say then that the characteristic factor for the averages in the left-hand side of the above is the Kronecker.
The proof of this theorem will follow easily from the following lemma.
@@
$\hskip 11.38092pt+\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F^{\prime}\rangle_{L^{2}(\mu\times\mu)}\langle F-F^{\prime},(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}$
$\hskip 11.38092pt+\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F-F^{\prime}\rangle_{L^{2}(\mu\times\mu)}\langle F,(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}$
@@
$\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}\big{|}\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F\rangle_{L^{2}(\mu\times\mu)}\big{|}^{2}=$
$=\sum_{1\leq i,j\leq k}\overline{c_{i}}c_{j}\|f_{2,i}^{\prime}\|_{2}\|f_{1,j}^{\prime}\|_{2}\|f_{2,j}^{\prime}\|_{2}\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}|\langle T_{g}f_{1},f_{1,i}^{\prime}\rangle|+\varepsilon=\varepsilon.$
@@
$+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}(f_{1}-\mathbb{E}_{\mu}(f_{1}\>|\>Z))\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z)$
$+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}\>|\>Z)\otimes T_{g}(f_{2}-\mathbb{E}_{\mu}(f_{2}\>|\>Z))$
(2.4) $+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}|Z)\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z).$
@@
A correspondence principle: Finally, we need the following instance of the correspondence principle, whose proof is classical, and we include it here for sake of completeness.
Theorem 2.14 (Correspondence principle).
@@
Consider the compact metric space $X:=\{0,1\}^{G}=\{x=(x_{g})_{g\in G}:x_{g}\in\{0,1\}\>\forall\>g\in G\}$, equipped with the Borel $\sigma$-algebra. We define the continuous action $T$ on $X$ by $T_{h}(x_{g})_{g\in G}=(x_{h^{-1}g})_{g\in G}$, for any $h\in G$ and $(x_{g})_{g\in G}\in X$. Now, consider the point $a=(\mathbbm{1}_{A}(g))_{g\in G}\in X$ and the clopen set $E=\{x=(x_{g})_{g\in G}\in X:x_{e_{G}}=1\}.$ By the choice of $a$, for $h\in G$ we have that $h\in A$ if and only if $T_{h}a\in E$, and therefore $A=\{h\in G:T_{h}a\in E\}.$ Consider the sequence of Borel probability measures on $X$ defined by
$N\mapsto\mu_{N}=\frac{1}{|\Phi_{N}|}\sum_{h\in\Phi_{N}}\delta_{T_{h}a}$
let $\mu^{\prime}$ be a ${}^{\ast}$ limit point of that sequence. Then $\mu^{\prime}(E)=d_{\Phi}(A)$, and $\mu^{\prime}$ is $T$-invariant, but not necessarily ergodic. Let $x\mapsto\mu_{x}^{\prime}$ be the ergodic decomposition of $\mu^{\prime}$. Then $\mu^{\prime}=\int_{x}\mu^{\prime}_{x}~{}\mathrm{d}\mu^{\prime}(x)$, so there exists $x_{0}\in X$ such that for the measure $\mu=\mu_{x_{0}}^{\prime}$, we have that $(X,\mu,T)$ is ergodic and $\mu(E)\geq~{}\mathrm{d}_{\Phi}(A)$. For all $N\in\mathbb{N}$, $\mu_{N}$ is supported on the orbit closure of $a$, and hence $\mu^{\prime}$ is also supported on the orbit closure of $a$. Therefore, we may assume that $\mu$ is also supported on the orbit closure of $a$ (as this is the case with $\mu^{\prime}_{x}$ for $\mu^{\prime}$-almost every $x\in X$). Then it follows by Lemma 2.9, that $a\in\texttt{{gen}}(\mu,\Psi)$ for some Følner sequence $\Psi$. ∎
@@ 3.1. Dynamical reformulation via correspondence principle
Usually in ergodic theory, correspondence principles serve as bridges between combinatorial and dynamical statements. Thus, we can use the correspondence principle (Theorem 2.14) to show that Theorem  1.5 follows from Theorem 3.1 below, which is more dynamical in nature.
@@ 3.2. Erdős progressions
The conclusion of Theorem 3.1 is still a rather combinatorial statement, so we need to reformulate it again into a dynamical statement. For this to be achieved, we will use the natural notion of Erdős progressions, as defined in [KMRR23b], which is a dynamical variant of arithmetic progressions.
@@
By assumption there exists an infinite sequence $(g_{n})_{n\in\mathbb{N}}$ in $G$ such that $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})$. Since $T_{g_{n}}a\to x_{1}\in E$ and $E$ is open, we get that $T_{g_{n}}a\in E$ for $n$ sufficiently large, so we may assume without loss of generality that $(g_{n})\subset\{g\in G:T_{g}a\in E\}$. Therefore we will construct the sequence $B$ to be a subset of $(g_{n})$.
@@
  • Since $T_{g_{n}}x_{1}\to x_{2}\in F$ and $F$ is open, we can pick $b_{1}\in(g_{n})$ such that $T_{b_{1}}x_{1}\in F$. Then $(x_{1},x_{2})\in(T^{-1}_{b_{1}}F)\times F$ which is open.
  • Since $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})$, we pick $b_{2}\in(g_{n})$ such that $(T_{b_{2}}\times T_{b_{2}})(a,x_{1})\in(T^{-1}_{b_{1}}F)\times F$ and $b_{2}\neq b_{1}$ (this is possible since there are infinitely many choices). Then
@@
  • Induction step: Assume we have found $b_{1},b_{2},\dots,b_{n}\in(g_{n})$ all distinct to each other such that
@@
    Since $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})\in\Big{(}\bigcap_{1\leq m\leq n}T^{-1}_{b_{m}}F\Big{)}\times F$ and this set is open, we can pick $b_{n+1}\in(g_{n})$ such that
@@ 3.3. Continuous factor maps to the Kronecker factor
We begin by generalizing a result of Host and Kra in [HK09] from actions of $(\mathbb{N},+)$ to actions of amenable groups.
@@
[HK09, Proposition 6.1 for group actions] Let $G$ be an amenable group, let $(X,\mu,T)$ be an ergodic $G$-system, $(Z,m,R)$ be its Kronecker factor and $\rho:(X,\mu,T)\to(Z,m,R)$ be a factor map. If $a\in X$ is a transitive point, then there exists a point $z\in Z$ and a Følner sequence $\Psi$ such that
@@
Let $(X,\mu,T)$ be an ergodic $G$-system, and let $a\in\texttt{{gen}}(\mu,\Phi)$ for some Følner sequence $\Phi$. Then there exists an ergodic extension $(\widetilde{X},\widetilde{\mu},\widetilde{T})$ of $(X,\mu,T)$, a Følner sequence $\widetilde{\Phi}$ and a point $\widetilde{a}\in\texttt{{gen}}(\widetilde{\mu},\widetilde{\Phi})$ such that:
@@
The proofs of (i) and (ii) are identical to those in the case $G$ is the semigroup $(\mathbb{N},+)$, and they can be found in [KMRR22, Proposition 3.20]. Therefore, here we only provide a sketch of the proof.
@@
Now, let us prove (iii). By assumption, $(\widetilde{T}_{g_{n}}\times\widetilde{T}_{g_{n}})(\widetilde{a},\widetilde{x}_{1})\to(\widetilde{x}_{1},\widetilde{x}_{2})$ for some $(g_{n})_{n\in\mathbb{N}}$ in $G$. We now notice that
@@ 4.1. Measures on Erdős progressions
Moreover, $m$ is the (left) Haar measure on $Z$, which is given as the push forward of the (left) Haar measure $m_{K}$ in $K$ by the natural projection $K\to Z=K/H$. We remark that since $K$ is compact, it is unimodular, so $m_{K}$, and consequently $m$, are two-sided invariant. Finally, the action $R=(R_{g})_{g\in G}$ is given by $R_{g}(z)=\alpha(g)z$, where $\alpha:G\to K$ is a continuous group homomorphism with dense image. Also, $\pi_{i}:X\times X\to X$ denotes the projection to the $i$-th coordinate, for $i=1,2$. Moreover, $z\mapsto\eta_{z}$ is a fixed disintegration of $\mu$ over the continuous factor map $\pi$.
@@
converges in the ${}^{\ast}$ topology to the Haar measure $m_{K}$ on $K$.
@@
The space of Borel probability measures on $K$ is ${}^{\ast}$ compact and metrizable, so in order to prove the result, it suffices to prove that if $(\nu_{N_{j}})_{j\in\mathbb{N}}$ is a convergent subsequence of $(\nu_{N})_{N\in\mathbb{N}}$, then it converges to the Haar measure $m_{K}$.
Let $(\nu_{N_{j}})_{j\in\mathbb{N}}$ be a subsequence of $(\nu_{N})_{N\in\mathbb{N}}$ which converges in the ${}^{\ast}$ topology to a measure $\nu$ on $K$. To prove that $\nu$ is the Haar measure on $K$, it suffices to prove that for all continuous functions $h$ on $K$ and all $k\in K$, we have that $\int_{K}h(ky)~{}\mathrm{d}\nu(y)=\int_{K}h(y)~{}\mathrm{d}\nu(y)$.
@@
$\bigg{|}\int_{K}h(y)~{}\mathrm{d}\nu(y)-\int_{K}h(ky)~{}\mathrm{d}\nu(y)\bigg{|}$ $\leq\bigg{|}\int_{K}h(y)~{}\mathrm{d}\nu(y)-\int_{K}h(\alpha(g_{0})y)~{}\mathrm{d}\nu(y)\bigg{|}+$
$\>\>\>\>\>\>\bigg{|}\int_{K}h(\alpha(g_{0})y)~{}\mathrm{d}\nu(y)-\int_{K}h(ky)~{}\mathrm{d}\nu(y)\bigg{|}$
@@
$m_{K}(A)=\sup_{\begin{subarray}{c}C\subset A\\ C\>\text{compact}\end{subarray}}m_{K}(C)=\inf_{\begin{subarray}{c}O\subset A\\ O\>\text{open}\end{subarray}}m_{K}(O)\>\>\text{ and }\>\>m_{K^{2}}(A)=\sup_{\begin{subarray}{c}C\subset A\\ C\>\text{compact}\end{subarray}}m_{K^{2}}(C)=\inf_{\begin{subarray}{c}O\subset A\\ O\>\text{open}\end{subarray}}m_{K^{2}}(O).$
@@
$0=m_{K}(C)=\inf_{\begin{subarray}{c}O\subset A\\ O\>\text{open}\end{subarray}}m_{K}(O),$
@@
Now, we define measures $\sigma$ on $X\times X$, and we want $\sigma$ to be defined as natural measures to study Erdős progressions. It is not hard to see that Erdős progressions on $K/H$ are exactly the triplets of the form $(z,kz,k^{2}z)$ for some $k\in K$ and $z\in K/H$. Therefore, following the definition given in [KMRR23b], we will define these measures as the natural measures on points $(x_{1},x_{2})\in X\times X$ to find Erdős progressions on $K/H$ starting at $\pi(a)$, namely $(\pi(a),k\pi(a),k^{2}\pi(a))$, for $k\in K$.
@@
Let us comment on why the measure $\sigma$ is indeed well-defined. Let $k_{0}\in K$ such that $\pi(a)=k_{0}H$. Since all the maps involved are measurable, we get that the map $K\to M(X)$, $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}~{}\mathrm{d}m_{K}(k)$ is indeed a Borel measurable map on $K$. It remains to prove that it is $m_{K}$ almost everywhere defined, and then the integral will induce a well defined Borel probability measure on $X\times X$.
Since $\eta$ is $m$-almost everywhere well-defined on $Z$, we can consider a set $Z^{\prime}\subset Z$ with $m(Z^{\prime})=1$ such that $\eta_{z}$ is well-defined for all $z\in Z^{\prime}$. Since $pm_{K}=m$, the set $K^{\prime}=p^{-1}(Z^{\prime})$ has $m_{K}(K^{\prime})=1$. Then also $m_{K}(K^{\prime}k_{0}^{-1})=1$, and then it is not difficult to check that the map $K\to M(X)$, $k\mapsto\eta_{k\pi(a)}$ is well defined on $K^{\prime}k_{0}^{-1}$. On the other hand, from Lemma 4.2, we have that $m_{K^{2}}$ is absolutely continuous with respect to $m_{K}$, so $1=m_{K^{2}}(K^{\prime}k_{0}^{-1})=m_{K}(s_{K}^{-1}(K^{\prime}k_{0}^{-1}))$, and again it is not too difficult to check that the map $K\to M(X)$, $k\mapsto\eta_{k^{2}\pi(a)}$ is well defined on $s_{K}^{-1}(K^{\prime}k_{0}^{-1})$.
So, after all, the map $K\to M(X\times X)$, $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}$ is well defined on $K^{\prime}k_{0}^{-1}\cap s_{K}^{-1}(K^{\prime}k_{0}^{-1})$ and $m_{K}(K^{\prime}k_{0}^{-1}\cap s_{K}^{-1}(K^{\prime}k_{0}^{-1}))=1$, i.e., $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}$ is $m_{K}$-almost everywhere well-defined, which was to be proved, and therefore $\sigma$ is indeed well-defined.
@@
Let $E\subset X$ with $\mu(E)>0$ and recall that we want to show that the set $E\times\bigcup_{t\in G}T^{-1}_{t}E$ has positive measure $\sigma$. We begin by expressing this set as

### v3

@@ 2. Preliminaries
[17, Theorem 1] Let $(X,\mu,T)$ be an ergodic $G$-system. Then its Kronecker factor is measurably isomorphic to a rotation on some homogeneous space $Z$ by some $\alpha$ with dense image.
From Proposition 2.12 we get that if $\mathcal{C}$ is the Borel $\sigma$-algebra on $Z$, then $\pi^{-1}(\mathcal{C})$ is equivalent to $\mathcal{A}$, i.e., they are equal modulo sets that have zero $\mu$ measure. Proposition 2.12 allows us to identify the Kronecker factor with a rotation on a homogeneous space, whenever $(X,\mu,T)$ is ergodic.
Characteristic factors for $G$-systems: The notion of characteristic factors will play a fundamental role later in one of our proofs. Here we have the following theorem for the characteristic factors with respect to some double averages that will concern us.
@@
(2.3) $\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}f_{1}\otimes T_{g}f_{2}=\lim_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}\>|\>Z)\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z)$
in $L^{2}(X\times X,\mu\times\mu)$.
Theorem 2.13 says that the Kronecker factor is the characteristic factor for the averages in the left-hand side of (2.3). The proof of Theorem 2.13 will follow easily from the next lemma.
@@
$\hskip 126.61467pt+\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F^{\prime}\rangle_{L^{2}(\mu\times\mu)}\langle F-F^{\prime},(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}$
$\hskip 126.61467pt+\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F-F^{\prime}\rangle_{L^{2}(\mu\times\mu)}\langle F,(T_{g}\times T_{g})(f_{1}\otimes f_{2})\rangle_{L^{2}(\mu\times\mu)}$
@@
$\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}\big{|}\langle(T_{g}\times T_{g})(f_{1}\otimes f_{2}),F\rangle_{L^{2}(\mu\times\mu)}\big{|}^{2}\\ =\sum_{1\leq i,j\leq k}\overline{c_{i}}c_{j}\|f_{2,i}^{\prime}\|_{2}\|f_{1,j}^{\prime}\|_{2}\|f_{2,j}^{\prime}\|_{2}\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}|\langle T_{g}f_{1},f_{1,i}^{\prime}\rangle|+\varepsilon=\varepsilon.$
@@
$\hskip 1.13791pt+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}(f_{1}-\mathbb{E}_{\mu}(f_{1}\>|\>Z))\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z)$
$\hskip 1.13791pt+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}\>|\>Z)\otimes T_{g}(f_{2}-\mathbb{E}_{\mu}(f_{2}\>|\>Z))$
(2.5) $\hskip 1.13791pt+\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}T_{g}\mathbb{E}_{\mu}(f_{1}|Z)\otimes T_{g}\mathbb{E}_{\mu}(f_{2}\>|\>Z).$
@@
A correspondence principle: Finally, we need the following instance of Furstenberg’s correspondence principle.
Theorem 2.15 (Cf. [4, Theorem 2.8]).
@@
Consider the compact metric space $X:=\{0,1\}^{G}=\{x=(x_{g})_{g\in G}:x_{g}\in\{0,1\}\>\forall\>g\in G\}$, equipped with the Borel $\sigma$-algebra. We define the continuous action $T$ on $X$ by $T_{h}(x_{g})_{g\in G}=(x_{gh})_{g\in G}$, for any $h\in G$ and $(x_{g})_{g\in G}\in X$. Now, consider the point $a=(\mathbbm{1}_{A}(g))_{g\in G}\in X$ and the clopen set $E=\{x=(x_{g})_{g\in G}\in X:x_{e_{G}}=1\}.$ By the choice of $a$, for $h\in G$ we have that $h\in A$ if and only if $T_{h}a\in E$, and therefore $A=\{h\in G:T_{h}a\in E\}.$ Consider the sequence of Borel probability measures on $X$ defined by
$N\mapsto\mu_{N}=\frac{1}{|\Phi_{N}|}\sum_{h\in\Phi_{N}}\delta_{T_{h}a},$
and let $\mu^{\prime}$ be a weak limit point of that sequence. Then $\mu^{\prime}(E)=d_{\Phi}(A)$, and $\mu^{\prime}$ is $T$-invariant, but not necessarily ergodic. Let $x\mapsto\mu_{x}^{\prime}$ be the ergodic decomposition of $\mu^{\prime}$. Then $\mu^{\prime}=\int_{x}\mu^{\prime}_{x}~{}\mathrm{d}\mu^{\prime}(x)$, so there exists $x_{0}\in X$ such that for the measure $\mu=\mu_{x_{0}}^{\prime}$, we have that $(X,\mu,T)$ is ergodic and $\mu(E)\geq~{}\mathrm{d}_{\Phi}(A)$. For all $N\in\mathbb{N}$, $\mu_{N}$ gives full measure to the orbit closure of $a$, and hence $\mu^{\prime}$ also gives full measure to the orbit closure of $a$. Therefore, we may assume that $\mu$ is also supported on the orbit closure of $a$ (as this is the case with $\mu^{\prime}_{x}$ for $\mu^{\prime}$-almost every $x\in X$). Then it follows by Lemma 2.9, that $a\in\texttt{{gen}}(\mu,\Psi)$ for some Følner sequence $\Psi$. ∎
@@ 3.1. Dynamical reformulation via correspondence principle
Usually in ergodic theory, correspondence principles serve as bridges between combinatorial and dynamical statements. In our situation, we can use the correspondence principle (Theorem 2.15) to show that Theorem  1.6 follows from Theorem 3.1 below, which is more dynamical in nature.
@@ 3.2. Erdős progressions
The conclusion of Theorem 3.1 is still a rather combinatorial statement, so we need to reformulate it again into a dynamical statement. For this to be achieved, we will use the notion of Erdős progressions, as defined in [15], which in $\mathbb{Z}$ is a dynamical variant of $3$-term arithmetic progressions. In our case, Erdős progressions are a dynamical variant of progressions of the form $(z,kz,k^{2}z)$, $k\in K$, $z\in Z$, which are the natural generalization of $3$-term arithmetic progressions in our setting.
@@
By assumption there exists an infinite sequence $(g_{n})_{n\in\mathbb{N}}$ in $G$ such that $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})$. Since $T_{g_{n}}a\to x_{1}\in E$ and $E$ is open, we get that $T_{g_{n}}a\in E$ for $n$ sufficiently large, so we may assume without loss of generality that $(g_{n})_{n\in\mathbb{N}}\subset\{g\in G:T_{g}a\in E\}$. Therefore we will construct the sequence $B$ to be a subset of $(g_{n})_{n\in\mathbb{N}}$.
@@
  • Since $T_{g_{n}}x_{1}\to x_{2}\in F$ and $F$ is open, we can pick $b_{1}\in(g_{n})_{n\in\mathbb{N}}$ such that $T_{b_{1}}x_{1}\in F$. Then $(x_{1},x_{2})\in(T^{-1}_{b_{1}}F)\times F$ which is open.
  • Since $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})$, we pick $b_{2}\in(g_{n})_{n\in\mathbb{N}}$ such that $(T_{b_{2}}\times T_{b_{2}})(a,x_{1})\in(T^{-1}_{b_{1}}F)\times F$ and $b_{2}\neq b_{1}$ (this is possible since there are infinitely many choices). Then
@@
  • Induction step: Assume we have found $b_{1},b_{2},\dots,b_{n}\in(g_{n})_{n\in\mathbb{N}}$ all distinct to each other such that
@@
    Since $(T_{g_{n}}\times T_{g_{n}})(a,x_{1})\to(x_{1},x_{2})\in\Big{(}\bigcap_{1\leq m\leq n}T^{-1}_{b_{m}}F\Big{)}\times F$ and this set is open, we can pick $b_{n+1}\in(g_{n})_{n\in\mathbb{N}}$ such that
@@ 3.3. Continuous factor maps to the Kronecker factor
We begin by generalizing a result of Host and Kra in [9] from actions of $(\mathbb{N},+)$ to actions of amenable groups.
@@
[9, Proposition 6.1 for group actions] Let $G$ be an amenable group, let $(X,\mu,T)$ be an ergodic $G$-system, $(Z,m,R)$ be its Kronecker factor and $\rho:(X,\mu,T)\to(Z,m,R)$ be a factor map. If $a\in X$ is a transitive point, then there exists a point $z\in Z$ and a Følner sequence $\Psi$ such that
@@
Let $(X,\mu,T)$ be an ergodic $G$-system, let $a\in\texttt{{gen}}(\mu,\Phi)$ for some Følner sequence $\Phi$, and $E\subset X$ be a set with $\mu(E)>0$. Then there exists an ergodic extension $(\widetilde{X},\widetilde{\mu},\widetilde{T})$ of $(X,\mu,T)$, a Følner sequence $\widetilde{\Phi}$ and a point $\widetilde{a}\in\texttt{{gen}}(\widetilde{\mu},\widetilde{\Phi})$ such that:
@@
The proofs of (i) and (ii) are identical to those in the case $G$ is the semigroup $(\mathbb{N},+)$, and they can be found in [14, Proposition 3.20]. Therefore, here we only provide a sketch of the proof.
@@
Note that (iii) is immediate from the definition of factors and factor maps. Hence, it remains to prove (iv). By assumption, $(\widetilde{T}_{g_{n}}\times\widetilde{T}_{g_{n}})(\widetilde{a},\widetilde{x}_{1})\to(\widetilde{x}_{1},\widetilde{x}_{2})$ for some $(g_{n})_{n\in\mathbb{N}}$ in $G$. We now notice that
@@ 4.1. Measures on Erdős progressions
Moreover, $m$ is the (left) Haar measure on $Z$, which is given as the push forward of the (left) Haar measure $m_{K}$ in $K$ by the natural projection $K\to Z=K/H$. We remark that since $K$ is compact, it is unimodular, so $m_{K}$ is two-sided invariant. Finally, the action $R=(R_{g})_{g\in G}$ is given by $R_{g}(z)=\alpha(g)z$, where $\alpha:G\to K$ is a group homomorphism with dense image. Also, $\pi_{i}:X\times X\to X$ denotes the projection to the $i$-th coordinate, for $i=1,2$. Moreover, $z\mapsto\eta_{z}$ is a fixed disintegration of $\mu$ over the continuous factor map $\pi$.
@@
converges in the weak topology to the Haar measure $m_{K}$ on $K$.
@@
The space of Borel probability measures on $K$ is weak compact and metrizable, so in order to prove the result, it suffices to prove that if $(\nu_{N_{j}})_{j\in\mathbb{N}}$ is a convergent subsequence of $(\nu_{N})_{N\in\mathbb{N}}$, then it converges to the Haar measure $m_{K}$.
Let $(\nu_{N_{j}})_{j\in\mathbb{N}}$ be a subsequence of $(\nu_{N})_{N\in\mathbb{N}}$ which converges in the weak topology to a measure $\nu$ on $K$. To prove that $\nu$ is the Haar measure on $K$, it suffices to prove that for all continuous functions $h$ on $K$ and all $k\in K$, we have that $\int_{K}h(ky)~{}\mathrm{d}\nu(y)=\int_{K}h(y)~{}\mathrm{d}\nu(y)$.
@@
$\bigg{|}\int_{K}h(y)~{}\mathrm{d}\nu(y)-\int_{K}h(ky)~{}\mathrm{d}\nu(y)\bigg{|}$ $\leq\bigg{|}\int_{K}h(y)~{}\mathrm{d}\nu(y)-\int_{K}h(\alpha(g_{0})y)~{}\mathrm{d}\nu(y)\bigg{|}$
$\hskip 1.13791pt+\bigg{|}\int_{K}h(\alpha(g_{0})y)~{}\mathrm{d}\nu(y)-\int_{K}h(ky)~{}\mathrm{d}\nu(y)\bigg{|}$
@@
$m_{K}(A)=\sup_{\begin{subarray}{c}C\subset A\\ C\>\text{compact}\end{subarray}}m_{K}(C)=\inf_{\begin{subarray}{c}O\supset A\\ O\>\text{open}\end{subarray}}m_{K}(O)\>\>\text{ and }\>\>m_{K^{2}}(A)=\sup_{\begin{subarray}{c}C\subset A\\ C\>\text{compact}\end{subarray}}m_{K^{2}}(C)=\inf_{\begin{subarray}{c}O\supset A\\ O\>\text{open}\end{subarray}}m_{K^{2}}(O).$
@@
$0=m_{K}(C)=\inf_{\begin{subarray}{c}O\supset C\\ O\>\text{open}\end{subarray}}m_{K}(O),$
@@
Now, we define a measure $\sigma$ on $X\times X$, and we want $\sigma$ to be defined as a natural measure to study Erdős progressions. It is not hard to see that Erdős progressions on $K/H$ are exactly the triplets of the form $(z,kz,k^{2}z)$ for some $k\in K$ and $z\in K/H$. Therefore, following the definition given in [15], we will define these measures as the natural measures on points $(x_{1},x_{2})\in X\times X$ to find Erdős progressions on $K/H$ starting at $\pi(a)$, namely $(\pi(a),k\pi(a),k^{2}\pi(a))$, for $k\in K$.
@@
Let us comment on why the measure $\sigma$ is well-defined. Let $k_{0}\in K$ be such that $\pi(a)=k_{0}H$. Since $\eta$ is $m$-almost everywhere defined on $Z$ and Borel measurable, we can consider a Borel measurable set $Z^{\prime}\subset Z$ with $m(Z^{\prime})=1$ such that $\eta_{z}$ is defined for all $z\in Z^{\prime}$. Since $p$ is a Borel measurable map and $pm_{K}=m$, the set $K^{\prime}=p^{-1}(Z^{\prime})$ is Borel measurable and has $m_{K}(K^{\prime})=1$. Then also $m_{K}(K^{\prime}k_{0}^{-1})=1$, and then it is not difficult to check that the map $K\to M(X)$, $k\mapsto\eta_{k\pi(a)}$ is defined on $K^{\prime}k_{0}^{-1}$. On the other hand, from Lemma 4.2, we have that $m_{K^{2}}$ is absolutely continuous with respect to $m_{K}$, so $1=m_{K^{2}}(K^{\prime}k_{0}^{-1})=m_{K}(s_{K}^{-1}(K^{\prime}k_{0}^{-1}))$, and again it is not too difficult to check that the map $K\to M(X)$, $k\mapsto\eta_{k^{2}\pi(a)}$ is defined on $s_{K}^{-1}(K^{\prime}k_{0}^{-1})$. So, after all, the map $K\to M(X\times X)$, $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}$ is defined on $K^{\prime}k_{0}^{-1}\cap s_{K}^{-1}(K^{\prime}k_{0}^{-1})$ and $m_{K}(K^{\prime}k_{0}^{-1}\cap s_{K}^{-1}(K^{\prime}k_{0}^{-1}))=1$, i.e., $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}$ is $m_{K}$-almost everywhere defined. Also, since all the maps involved in the definition of $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}$ are Borel measurable, we have that $k\mapsto\eta_{k\pi(a)}\times\eta_{k^{2}\pi(a)}$ is also Borel measurable. Therefore, $\sigma$ is indeed well-defined.
@@
Let $E\subset X$ with $\mu(E)>0$ and recall that we want to show that the set $E\times\bigcup_{t\in G}T^{-1}_{t}E$ has positive measure with respect to $\sigma$. We begin by expressing this set as